OpenStudy (anonymous):

8=(square root) 2^(4x+8)

OpenStudy (anonymous):

8 = 16^(x+2) sqrt(x)

OpenStudy (anonymous):

\[8=2^3=\sqrt{2}^6\] so you can write \[6=4x+8\] \[4x=-2\] \[x=-\frac{1}{2}\]

OpenStudy (anonymous):

how can i explain that in sentences